OpenStudy (anonymous):

Can someone show me the steps to simplify this algebra problem please? $\frac{5+\sqrt{3}}{2-\sqrt{3}}$

jimthompson5910 (jim_thompson5910):

hint: multiply top and bottom by \(\large 2+\sqrt{3}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would it be \[10+\sqrt{9}\] after multipling?

jimthompson5910 (jim_thompson5910):

I'm assuming you got that after multiplying \(\large (5+\sqrt{3})(2+\sqrt{3})\)

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

but that's not correct since you didn't FOIL

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

hint: (a+b)(c+d) is the same as a(c+d) + b(c+d)

OpenStudy (anonymous):

I learned how to do this before but have forgotten how to do this type of math. I'm trying to help my nephew with his homework.

jimthompson5910 (jim_thompson5910):

FOIL stands for First, Outer, Inner, Last and it helps you multiply two binomials

jimthompson5910 (jim_thompson5910):

This page explains it better http://www.mathwarehouse.com/algebra/polynomial/foil-method-binomials.php

OpenStudy (anonymous):

when you multiple the \[\sqrt{3} and the \sqrt{3}\] does it equal \[\sqrt{9}\] and then simplify further down to just 3?

OpenStudy (anonymous):

I figured it out thank you

OpenStudy (anonymous):

\[13+7\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

that's correct and you're welcome